# VISAGE: Video Instance Segmentation with Appearance-Guided Enhancement

### 08 Dec 2023 v1

@@ Abstract
In recent years, online Video Instance Segmentation (VIS) methods have shown remarkable advancement with their powerful query-based detectors. Utilizing the output queries of the detector at the frame level, these methods achieve high accuracy on challenging benchmarks. However, we observe the heavy reliance of these methods on the location information that leads to incorrect matching when positional cues are insufficient for resolving ambiguities. Addressing this issue, we present  VISAGE that enhances instance association by explicitly leveraging appearance information. Our method involves a generation of queries that embed appearances from backbone feature maps, which in turn get used in our suggested simple tracker for robust associations. Finally, enabling accurate matching in complex scenarios by resolving the issue of over-reliance on location information, we achieve competitive performance on multiple VIS benchmarks. For instance, on YTVIS19 and YTVIS21, our method achieves 54.5 AP and 50.8 AP. Furthermore, to highlight appearance-awareness not fully addressed by existing benchmarks, we generate a synthetic dataset where our method outperforms others significantly by leveraging the appearance cue. Code will be made available at https://github.com/KimHanjung/VISAGE.
@@ 1 Introduction
As detectors directly impact the accuracy in the video domain, recent online models are primarily built using the powerful query-based detectors [38, 5]. Spatially decoding image information, the detectors are designed to represent object-wise information using the queries. Therefore, the online VIS methods reuse these queries from the detectors and have achieved substantial improvements in multiple challenging benchmarks [25, 33] by mostly adopting either propagation [16, 23] or matching [17, 29, 36] strategies. However, tracking under complex scenarios such as shot changes or trajectory intersections (Fig. 1) are yet imperfect, resulting in the degradation of the overall accuracy.
@@
Examining these failure cases, we observe that object-wise information of the queries are significantly imbalanced: heavy reliance on positional cues, and less reflection of appearances. As demonstrated in Fig. 1, previous query-based methods [16, 36, 17, 29, 37], display a tendency to maintain the spatial order of previous predictions in their current predictions. To support this argument, we conduct additional analyses by horizontally flipping images to generate two-frame pseudo videos. The existing models manifest association errors despite the distinct exterior patterns of objects (Fig. 2 Top), which highlight the dependence on object locations. As there exist multiple scenarios that cannot be fully handled with the imbalanced information, such a phenomenon necessitates the models to take object appearances into consideration.
We introduce VISAGE (Video Instance Segmentation with Appearance-Guided Enhancement), a method that leverages appearance cues as a crucial indicator for distinguishing instances. First, we present a lightweight branch for generating appearance queries that encapsulate the visual characteristics of an instance. Specifically, given mask predictions of object queries from the detector, this branch explicitly embeds appearance by performing mask pooling operations to backbone feature maps. Then, our method separately uses the contrastive loss [26, 3] to pairs of appearance queries and object queries, which improve the association of instances across a video by alleviating the position-appearance imbalance.
Additionally, we present a simple tracker that employs appearance cues as an additional indicator and leverages temporal information. Previous methods [29, 36] employ multiple refinement steps for removing redundant mask proposals and adopt handcrafted threshold values to improve the accuracy. While these intricate tracker modules raise the tracking quality, they become dependent on multiple hyperparameters that are heuristically optimized for each dataset. In order to alleviate such dependence, our method streamlines the tracker and dramatically reduces the number of hyperparameters, eliminating thresholding and instead incorporating appearance cues. Nonetheless, a lack of temporal information still exists, inherent in query-matching online methods, as they are only aware of adjacent frames. We address this limitation by using a simple memory bank to facilitate temporal awareness.
Our approach VISAGE has many desirable properties. Enhanced by the appearance guidance, VISAGE achieves robust tracking in complex scenarios, as illustrated at Fig. 1 (a) and bottom of Fig. 2. Furthermore, with its simplified tracker that effectively utilizes the past history of both object and appearance queries, VISAGE demonstrates competitive performance across all benchmark datasets. We validate our methodology through extensive evaluations on the four standard VIS benchmarks: YouTubeVIS-19/21/22[31, 32, 33] and OVIS [25]. Unfortunately, these datasets are mostly composed of videos with general scenarios, thus not suitable for precise analyses of the effectiveness of leveraging appearance embeddings. Therefore, to validate our method’s effectiveness in leveraging appearance cues, we develop a large-scale pseudo dataset and conduct comparisons of its performance with other methods. Finally, VISAGE demonstrates superior performance over previous methods [16, 36] in the comparison.
@@ 2.1 Online Video Instance Segmentation
Contrary to traditional online approaches [31, 34, 11, 19], modern online methods utilize query-based detectors with an emphasis on query association strategy. Query-based online methods can be divided into two main strategies: query-matching and query-propagation.
Query-matching approaches focus on dynamically building tracklets in an online manner. For instance, MinVIS [17] utilizes Mask2Former [5] as its query-based detector trained exclusively at the image level. During inference, MinVIS independently predicts outputs at the frame level and then associates them through bipartite matching between frame queries. To enhance the distinctiveness of queries, IDOL [29] introduces contrastive learning, inspired by QDTrack [8]. They also adopt same inference process with QDTrack, additionally propose temporal weighted softmax with memory bank for leveraging temporal information. However, those memory bank is not used at training stage. CTVIS [36] further improves upon IDOL by removing the discrepancy between training and inference caused by the use of the memory bank. By utilizing the memory bank during training, CTVIS eliminates this discrepancy and incorporates historical instances. All these methods rely on the outputs of query-based detectors for their matching processes. In contrast, our approach VISAGE, introduces additional query to ensure more stable matching.
@@
Query-propagation methods utilize output queries from previous frames to detect objects in the current frame. InsPro [12] propagates their instance query and proposal for association and proposes duplicate removal strategy. CAROQ [6] extends Mask2Former [5] for application in VIS methods and propagates its output queries along with relative object queries. GenVIS [16] employs an offline VIS model [15], and adapts it to a query-propagation approach, notably omitting heuristic association strategies. This method reduces computational overhead by freezing the backbone network and query-based detector, thus effectively handling longer video sequences. Similarly, DVIS [37] introduces referring cross-attention, which is designed to capture the correlation between the current and previous frames using their queries. Yet, these methods are limited by their local matching strategy, which focuses on aligning tracklets with ground truths. This approach often leads to unstable tracking outcomes, as it does not consider the full video context. TCOVIS [23] addresses this issue by shifting from the local matching strategy of GenVIS to a global matching approach, thereby achieving more robust tracking across videos.
2.2 Offline Video Instance Segmentation
Previous offline VIS methods [1, 30] treat video clip input as a spatio-temporal volume. However, with the advent of query-based detectors, recent offline methods have shifted to treating videos through queries. Recent offline methods mostly leverage query-based detectors. VisTR [27] builds upon DETR [2] to create an end-to-end framework. However, the self-attention mechanism in VisTR is computationally heavy. To address this, IFC [18] introduces memory tokens that allow communication between frames, reducing the computational overhead. SeqFormer [28] restructures the decoder to process each frame separately, enhancing the model’s sensitivity to instance movement by capturing each location. Mask2Former-VIS [4] extends the Mask2Former [5] for video-level tasks, achieving notable performance improvements. TeViT [35] presents a novel backbone that improves temporal information processing, leveraging the capabilities of ViT [7]. VITA [15] effectively aggregates video information by learning the relationships between the frame-level queries.
3 Method
We propose a novel online VIS framework that leverages appearance information to reduce the heavy reliance on location information. Our newly introduced appearance-aware query guides the association procedure, ensuring correct matching. Finally, we introduce a simple inference method that aligns with our design principles, without heuristics.
3.1 Query-based Detector
In most recent query-based object detection frameworks, the architecture is typically segmented into three integral components: the backbone, the transformer encoder, and the transformer decoder. Initially, the backbone generates the low-level image feature maps, which encapsulate the essential visual information. Subsequent to this, the transformer encoder employs self-attention mechanisms to enhance the feature representation, as detailed in [2, 38]. Finally, in the transformer decoder, the identified objects are decoded into $N$ learnable queries.
Our method uses the well-established query-based detection framework [5], maintaining the original structure with the backbone, transformer encoder, and transformer decoder. Furthermore, we utilize the feature maps from the backbone to create the appearance query, and we use the output of the transformer decoder as the object query.
3.2 Appearance-Guided Enhancement
Before the emergence of query-based tracking approaches, video tracking methods [31, 8, 20] generate instance features from backbone feature maps. This is achieved using operations like RoIPool [10] and RoIAlign [14], followed by associating them across the video. In line with these traditional methods, we use average pooling to extract appearance feature from the backbone feature maps, guided by the predicted masks as shown in Fig. 3 (a). Therefore, in addition to the object queries already in use, we further employ the appearance query as an additional indicator. We transform both appearance queries and object queries, leading to the creation of appearance embeddings $\mathbf{e}_{a}\in\mathbb{R}^{N\times C}$ and object embeddings $\mathbf{e}_{i}\in\mathbb{R}^{N\times C}$, respectively.
@@ 3.2 Appearance-Guided Enhancement
To ensure more distinctive embeddings for better association, we apply contrastive loss to both object and appearance embeddings, similar to the approach used in previous methods [29, 8]. We sample key frame and reference frame, then calculate the loss for each set of embeddings separately. Consequently, our final loss is a weighted summation of original query-based detector’s loss and the contrastive loss for both the appearance embeddings and object embeddings.
@@
1:model $\mathcal{F}$, frames $\{x^{t}\}_{t=1}^{T}$
@@
11:     $\mathbf{s}\leftarrow\texttt{cos_sim}(\mathbf{e}_{i}^{t},\mathbf{m}_{i}^{t})+\texttt{cos_sim}(\mathbf{e}_{a}^{t},\mathbf{m}_{a}^{t})$
@@
13:end for
14:return $\mathcal{P}$
3.3 Inference with Appearance
As shown in Algorithm 1, our tracker employs a simple yet effective matching process that aligns the current object embeddings and appearance embeddings with their respective counterparts from the previous frame. We compute the similarity scores for each object-appearance embedding pair using Cosine Similarity, following a method similar to that in [17]. Then, employing the sum of object and appearance similarities, we utilize the Hungarian algorithm [21] (linear_sum_assignment in Algorithm 1) to achieve optimal assignment.
@@ 4 Experiments
Method Setting YouTube-VIS 2019 YouTube-VIS 2021
[…]
IFC [18] offline 41.2 65.1 44.6 42.3 49.6 35.2 55.9 37.7 32.6 42.9
Mask2Former-VIS [4] offline 46.4 68.0 50.0 - - 40.6 60.9 41.8 - -
MinVIS [17] online 47.4 69.0 52.1 45.7 55.7 44.2 66.0 48.1 39.2 51.7
IDOL [29] online 49.5 74.0 52.9 47.7 58.7 43.9 68.0 49.6 38.0 50.9
VITA [15] offline 49.8 72.6 54.5 49.4 61.0 45.7 67.4 49.5 40.9 53.6
GenVIS [16] online 50.0 71.5 54.6 49.5 59.7 47.1 67.5 51.5 41.6 54.7
GenVIS [16] offline 51.3 72.0 57.8 49.5 60.0 46.3 67.0 50.2 40.6 53.2
DVIS [37] online 51.2 73.8 57.1 47.2 59.3 46.4 68.4 49.6 39.7 53.5
DVIS [37] offline 52.6 76.5 58.2 47.4 60.4 47.4 71.0 51.6 39.9 55.2
TCOVIS [23] online 52.3 73.5 57.6 49.8 60.2 49.5 71.2 53.8 41.3 55.9
CTVIS [36] online 55.1 78.2 59.1 51.9 63.2 50.1 73.7 54.7 41.8 59.5
VISAGE online 54.5 77.2 60.3 50.9 61.7 50.8 72.9 55.6 43.3 58.8
Table 1: Comparisons on the YouTube-VIS 2019 and 2021 validation sets. Methods are denoted as online or offline, indicated by the text color. We highlight the best performance in bold and the second-best performance with an underline.
@@ 4.1 Datasets
OVIS stands out for its more complex and longer videos, featuring 25 categories and comprising 2,238 videos for training, 302 for validation, and 343 for testing. In comparison to YTVIS, it contains a higher number of instances per video, averaging 5.8, and a total of 296k masks. Additionally, the average length of videos in OVIS is approximately 12 seconds.
@@ 4.2 Implementation Details
We adopt the Mask2Former [5] as our query-based detector. All of our models are initialized with parameters pre-trained on COCO dataset [24] with ResNet-50 [13] backbone. We also adopt COCO joint training following previous works [28, 15, 16, 36, 23]. Our batch includes 16 videos. In our experimental setup, we set the weights for the contrastive losses applied to both the appearance and object embeddings at $2.0$. We follow our detector’s loss weight configurations as detailed in  [5] for other losses in our setup. The window size of memory bank $W$ is set to 5 during inference. Finally, our method is trained using 4 NVIDIA A6000 GPUs.
@@
MinVIS [17] 23.3 47.9 19.3 20.2 28.0
VITA [15] 32.6 53.9 39.3 30.3 42.6
GenVIS [16] 37.5 61.6 41.5 32.6 42.2
TCOVIS [23] 38.6 59.4 41.6 32.8 46.7
VISAGE 36.7 60.0 41.1 35.4 45.1
Table 2: Comparison on the YTVIS 2022 long videos set. Bold and underline highlight the highest and second-highest performances, respectively. For methods that are both online and semi-online, the online version is used.

### 08 Mar 2024 v2

@@ Abstract
In recent years, online Video Instance Segmentation (VIS) methods have shown remarkable advancement with their powerful query-based detectors. Utilizing the output queries of the detector at the frame-level, these methods achieve high accuracy on challenging benchmarks. However, our observations demonstrate that these methods heavily rely on location information, which often causes incorrect associations between objects. This paper presents that a key axis of object matching in trackers is appearance information, which becomes greatly instructive under conditions where positional cues are insufficient for distinguishing their identities. Therefore, we suggest a simple yet powerful extension to object decoders that explicitly extract embeddings from backbone features and drive queries to capture the appearances of objects, which greatly enhances instance association accuracy. Furthermore, recognizing the limitations of existing benchmarks in fully evaluating appearance awareness, we have constructed a synthetic dataset to rigorously validate our method. By effectively resolving the over-reliance on location information, we achieve state-of-the-art results on YouTube-VIS 2019/2021 and Occluded VIS (OVIS). Code is available at https://github.com/KimHanjung/VISAGE.
Keywords:
video instance segmentation, object appearance
@@ 1 Introduction
As detectors directly impact the accuracy in the video domain, recent online models are primarily built using the powerful query-based detectors [39, 5]. Spatially decoding image information, the detectors are designed to represent object-wise information using the queries. Therefore, the online VIS methods reuse these queries from the detectors and have achieved substantial improvements in multiple challenging benchmarks [26, 34] by mostly adopting either propagation [15, 23] or matching [17, 31, 37] strategies. However, tracking under complex scenarios such as shot changes or trajectory intersections (Fig. 1) remains imperfect, resulting in the degradation of the overall accuracy.
@@
Examining these failure cases, we observe that object-wise information of the queries is significantly imbalanced: heavy reliance on positional cues, and less reflection on appearances. As demonstrated in Fig. 1, previous query-based VIS methods [15, 37, 17, 31, 38], tend to maintain the spatial order of previous predictions in their current predictions. To support this argument, we conduct additional analyses by horizontally flipping images to generate two-frame pseudo videos. The existing models manifest association errors despite the distinct exterior patterns of objects, as shown at the top of Fig. 2, which highlight the dependence on object locations. As there exist multiple scenarios that cannot be fully handled with the imbalanced information, such a phenomenon necessitates the models to take object appearances into consideration.
We introduce VISAGE (Video Instance Segmentation with Appearance-Guided Enhancement), a method that leverages appearance cues as a crucial indicator for distinguishing instances. In our approach, we introduce a streamlined branch that employs mask pooling to generate appearance queries from the predicted mask of object queries. This enables each appearance query to capture the visual features of its corresponding object, providing a more comprehensive representation for improved tracking accuracy. To refine query discrimination, we integrate a contrastive loss [27, 3, 8], which enhances the model’s ability to distinguish between instances across different frames.
@@
Additionally, we introduce a streamlined tracker designed to minimize reliance on heuristic procedures to the greatest extent feasible. Previous methods [31, 37] employ multiple refinement steps for removing redundant mask proposals and adopt handcrafted threshold values for accurate tracklet construction. Through such processing, only selectively chosen queries are incorporated into the matching process. Although these complex tracker configurations enhance tracking accuracy, they also create a dependency on numerous hyperparameters, each of which can be tailored heuristically to specific datasets. To alleviate such dependence, our method streamlines the tracker and dramatically reduces the number of required hyperparameters, such as threshold values for initializing and deleting tracklets, and non-maximum suppression (NMS), among others. Nonetheless, a lack of temporal information still exists, inherent in query-matching online methods, as they are only aware of adjacent frames. We address this limitation by using a simple memory bank to facilitate temporal awareness.
Despite the simplicity of our approach, VISAGE has many desirable properties. Our method introduces a new paradigm in query-based VIS by emphasizing the crucial role of appearance information for object association. Enhanced by appearance-based guidance, it demonstrates superior performance in complex tracking scenarios, outperforming previous methods that often misidentify objects due to an excessive dependence on spatial information as shown in Fig. 1(a). It successfully leverages appearance information, as illustrated at the bottom of Fig. 2, and has been validated on our proposed large-scale pseudo dataset, outperforming established methods [15, 37] by a large margin. Furthermore, with its simplified tracker that effectively utilizes the past history of both object and appearance queries, our method demonstrates competitive performance across all benchmark datasets. Notably, VISAGE achieves state-of-the-art performance on three standard benchmarks: YouTubeVIS-19/21 [32, 33] and OVIS [26].
@@ 2.1 Online Video Instance Segmentation
Contrary to traditional online approaches [32, 35, 11, 19], modern online methods utilize query-based detectors [2, 39, 5] with an emphasis on query association strategies. Query-based online methods can be divided into two main strategies: query-matching and query-propagation.
Query-matching approaches dynamically construct tracklets in an online fashion using query-based detectors  [2, 39, 5], which yield predictions for each video frame individually. MinVIS [17] implements this concept by exclusively training of the query-based detector, subsequently deploying it on video frames independently during the inference phase to conduct tracking via bipartite matching of corresponding queries. However, its supervision is restricted to the frame-level, which can introduce ambiguity by not accounting for the object’s continuity in the video sequence. To mitigate this, some studies [31, 37] have incorporated contrastive learning [27, 3, 8] to refine instance embeddings. Moreover, these methods utilize a memory bank at the inference stage, which allows for the processing of multiple frames, thereby enhancing the temporal information captured for each object. Specifically, CTVIS [37] further improves discriminative capability by using the memory bank during the training stage. However, these previous approaches still fall short in discriminative ability due to their insufficient use of appearance information. Our method, VISAGE, introduces a novel strategy that utilizes the appearance cue to significantly enhance discriminative ability, leading to more robust tracking.
Query-propagation methods track objects in video sequences by utilizing output queries from prior frames. By propagating output queries, these methods track corresponding objects across frames [6]. Additionally, some enhance tracking accuracy by also propagating proposals alongside the queries [12]. Recent developments have enabled some methods to operate in both online and offline modes, processing videos clip-by-clip and frame-by-frame. For instance, GenVIS [15] employs an offline VIS method as its backbone, wherein learnable instance prototypes aggregate the backbone’s outputs through the propagation of the instance prototype. This design allows GenVIS to function as either an online or offline method, depending on the length of clip processed by the backbone. On the other hand, DVIS [38] constructs tracklets by propagating frame queries in an online manner. These well-established tracklets are then refined to effectively utilize information from the entire video in an offline manner. Yet, these methods are often constrained by a local matching strategy that focuses on aligning tracklets with ground truths based solely on the current frame. This approach frequently leads to unstable tracking outcomes, as it fails to consider the entire video context, resulting in suboptimal performance. TCOVIS [23] addresses this issue by shifting from the local matching strategy of GenVIS [15] to a global matching approach, thereby achieving more robust tracking across videos.
2.2 Offline Video Instance Segmentation
Offline VIS architectures [1, 29] process input videos at the clip-level rather than frame by frame. VisTR [28] extends DETR [2] from frame-level to clip-level processing in an end-to-end manner by simultaneously handling multiple frame features within the transformer encoder-decoder. However, this approach, which processes multiple frame-level inputs at once, demands extensive computation, making the processing of longer sequences impractical.
To overcome this limitation, IFC [18] introduces a memory token in the transformer encoder and employs fixed-size clip queries in the transformer decoder, enhancing both the model’s performance and efficiency. With the transformer decoder design of IFC, Mask2Former-VIS [4] adapts Mask2Former [5] for video-level tasks, resulting in substantial performance gains. Meanwhile, SeqFormer [30] redesigns the transformer decoder to process each frame individually, thus improving the model’s ability to detect instance movement by precisely capturing location changes. VITA [16] introduces a novel strategy by processing clip queries through cross-attention with frame queries rather than relying on frame features. This strategy lightens the computational load imposed by dense spatio-temporal backbones, resulting in an efficient architecture capable of managing lengthy videos. Beyond focusing solely on transformer architecture enhancements, TeViT [36] introduces a groundbreaking backbone that improves temporal information processing by capitalizing on the strengths of ViT [7].
3 Method
We now present VISAGE. Initially, we provide an overview of the query-based detector architecture upon which VISAGE is constructed. Then, we introduce a novel strategy that improves object association through appearance guidance. Lastly, we detail our inference pipeline with a simplified structure that significantly enhances performance.
3.1 Query-based Detector
Query-based object detectors [2, 39, 5] can be largely divided into three components: backbone, transformer encoder, and transformer decoder. The backbone initiates the process by generating low-level image feature maps, encapsulating essential visual information. These feature maps are then enhanced through the transformer encoder, which employs self-attention mechanisms to refine the feature representation, as detailed in [2, 39]. The process concludes in the transformer decoder, where the identified objects are decoded into $N$ learnable queries. Our approach adopts the well-established query-based detector framework [5], maintaining its original structure intact.
3.2 Appearance-Guided Enhancement
Prior to the advent of query-based tracking approaches, traditional video tracking methods [32, 8, 20] extracted instance features from backbone feature maps using operations such as RoIPool [10] and RoIAlign [13]. Following a similar principle, we use average pooling to extract appearance queries from the backbone feature maps, guided by the predicted masks as shown in Fig. 3 (a). These appearance queries are designed to encapsulate appearance-centric features, providing a distinctive complement to the object queries. Consequently, alongside the object queries already in use, we introduce appearance queries as an additional indicator. We then transform both types of queries into appearance embeddings $\mathbf{e}_{a}\in\mathbb{R}^{N\times C}$ and object embeddings $\mathbf{e}_{i}\in\mathbb{R}^{N\times C}$, respectively.
@@ 3.2 Appearance-Guided Enhancement
To enhance object association quality, we improve the distinctiveness of both object and appearance embeddings. We utilize contrastive learning to refine embeddings obtained from two distinct frames, ensuring that embeddings of identical object instances are brought closer together in the embedding space, while those of different instances are separated further apart. Unlike previous methods utilizing contrastive learning [31, 8], our approach treats object and appearance embeddings individually, applying contrastive loss to each respectively. It allows each type of embedding to be distinctly characterized by its inherent properties, facilitating their mutual synergy in object matching, as illustrated in Fig. 3 (c). Consequently, our model’s final loss integrates a weighted sum of the contrastive losses for both types of embeddings with the original query-based detector’s loss.
3.3 Inference with Appearance
As shown in Algorithm 1, our tracker employs a simple yet effective matching process that aligns the current object embeddings and appearance embeddings with their respective counterparts from the previous frame. We compute the similarity scores for each object-appearance embedding pair using Cosine Similarity, following a method similar to that in [17]. Then, employing the weighted sum of object and appearance similarities, we utilize the Hungarian algorithm [21] (linear_sum_assignment in Algorithm 1) to achieve optimal assignment.
@@ 3.3 Inference with Appearance
1:model $\mathcal{F}$, frames $\{x^{t}\}_{t=1}^{T}$, weight $\alpha$
@@
11:     $\mathbf{s}\leftarrow(1-\alpha)\cdot\texttt{cos_sim}(\mathbf{e}_{i}^{t},\mathbf{m}_{i}^{t})$
12:        $+\ \alpha\cdot\texttt{cos_sim}(\mathbf{e}_{a}^{t},\mathbf{m}_{a}^{t})$
@@
14:end for
15:return $\mathcal{P}$
@@ 4.1 Datasets
OVIS stands out for its more complex and longer videos, featuring 25 categories and comprising 607 videos for training, 140 for validation, and 154 for testing. In comparison to YTVIS, it contains a higher number of instances per video, averaging 5.8, and a total of 296k masks. Additionally, the average length of videos in OVIS is approximately 12 seconds.
@@ 4.2 Implementation Details
We adopt the Mask2Former [5] as our query-based detector. All of our models are initialized with parameters pre-trained on COCO dataset [24] with ResNet-50 [14] backbone. We also adopt COCO joint training following previous works [30, 16, 15, 37, 23]. Our batch includes 16 videos. In our experimental setup, we set the weights for the contrastive losses applied to both the appearance and object embeddings at $2.0$. For losses other than the contrastive loss, we adopt the same loss function and weight specifications as those described in [4]. The window size of memory bank $W$ is set to 5 and appearance weight $\alpha$ is set to 0.75 during inference. Finally, our method is trained using 4 NVIDIA A6000 GPUs.
Table 1: Comparisons on the YouTube-VIS 2019, 2021 and 2022 validation sets. Methods are denoted as online or offline, indicated by the text color. We highlight the best performance in bold.
Method Setting YouTube-VIS 2019 YouTube-VIS 2021 YouTube-VIS 2022
@@
IFC [18] offline 41.2 65.1 44.6 42.3 49.6 35.2 55.9 37.7 32.6 42.9 - - - - -
Mask2Former-VIS [4] offline 46.4 68.0 50.0 - - 40.6 60.9 41.8 - - - - - - -
MinVIS [17] online 47.4 69.0 52.1 45.7 55.7 44.2 66.0 48.1 39.2 51.7 23.3 47.9 19.3 20.2 28.0
IDOL [31] online 49.5 74.0 52.9 47.7 58.7 43.9 68.0 49.6 38.0 50.9 - - - - -
VITA [16] offline 49.8 72.6 54.5 49.4 61.0 45.7 67.4 49.5 40.9 53.6 32.6 53.9 39.3 30.3 42.6
GenVIS [15] online 50.0 71.5 54.6 49.5 59.7 47.1 67.5 51.5 41.6 54.7 37.5 61.6 41.5 32.6 42.2
GenVIS [15] offline 51.3 72.0 57.8 49.5 60.0 46.3 67.0 50.2 40.6 53.2 37.2 58.5 42.9 33.2 40.4
DVIS [38] online 51.2 73.8 57.1 47.2 59.3 46.4 68.4 49.6 39.7 53.5 - - - - -
DVIS [38] offline 52.6 76.5 58.2 47.4 60.4 47.4 71.0 51.6 39.9 55.2 - - - - -
TCOVIS [23] online 52.3 73.5 57.6 49.8 60.2 49.5 71.2 53.8 41.3 55.9 38.6 59.4 41.6 32.8 46.7
CTVIS [37] online 55.1 78.2 59.1 51.9 63.2 50.1 73.7 54.7 41.8 59.5 - - - - -
VISAGE online 55.1 78.1 60.6 51.0 62.3 51.6 73.8 56.1 43.6 59.3 37.5 60.0 37.1 35.2 44.1